OpenStudy (anonymous):

integral as 0 approaches 1 (x-4)/x^2-5x+6 dx

OpenStudy (turingtest):

factor the denominator and apply partial fractions

OpenStudy (anonymous):

can you show me how to do that because i've tried and keep getting the wrong answer

OpenStudy (turingtest):

how would you factor x^2-5x+6 ?

OpenStudy (anonymous):

(x-3)(x-2)

OpenStudy (anonymous):

i got up to integral from 0-1 6/x-2 - 5/x-3

OpenStudy (anonymous):

Can you find A and B for this equation to be true? \[\int\limits_{0}^{1}{\frac{x-4}{(x-2)(x-3)}}dx=\int\limits_{0}^{1}(\frac{A}{x-2}+\frac{B}{x-3})dx\]

OpenStudy (anonymous):

i done all this . i just dont know how to integrate the ln

OpenStudy (anonymous):

Can you write the integral?

OpenStudy (anonymous):

If you mean this:

OpenStudy (anonymous):

i got 6ln(x-2) - 5ln(x-3) } 1-0

OpenStudy (anonymous):

Same principle i guess. Or you can write 6ln(x-2) as 6lnx/ln2=(6/ln2)*lnx where int lnx have been done.

OpenStudy (anonymous):

i have no clue, can you help me out from top to bottom?

OpenStudy (anonymous):

Can you write the integral you want to evaluate in equation mode first?

OpenStudy (anonymous):